### v1

@@ Abstract
In this paper we introduce the $q$-differential operator for $q$-spinor variables. We establish the $q$-spinor chain rule , the new $q$-differential operator, the $q$-Dirac differential operators and the $q$-complex spinor integrals. We also define the $q$-spinor differential equation. The suggestions for further work at the end of the paper.
@@ 1 Introduction
Based on the work of Beretetskii et al., Lachieze-Rey, Gori et al., and Cartan, the spinor $\psi^{\alpha}$ is defined as a magnitude components $\alpha=1,2$ expressed as $\psi^{\alpha}=\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}$ and its complex conjugate $\dot{\varphi}$ in terms of the rotation matrices (see [2],[1],[9],[5] for more details). Based on the work of the previously mentioned authors, there are two types of operations on spinors, which are reflections and rotations. In group theory, the set of rotations described by the matrices with complex entries is group $SU(2)$, whose generators are the Pauli matrices, described in the work of Zettili [17]. With respect to rotation matrices, Gori et al. mention, in their work, the rotation matrices that originated the Pauli matrices in the form:
$R_{x}(\theta)=\begin{bmatrix}\cos(\theta/2)&i\sin(\theta/2)\\ i\sin(\theta/2)&\cos(\theta/2)\end{bmatrix},R_{y}(\theta)=\begin{bmatrix}\cos(\theta/2)&\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\end{bmatrix},R_{z}(\theta)=\begin{bmatrix}e^{i\theta/2}&0\\ 0&e^{-i\theta/2}\end{bmatrix},$ (1)
being $\theta$ the angle of rotation [5]. Beretetskii et al. define the covariance and contravariance over the spinors by the relation $\psi^{\prime 1}=\psi_{2},\psi^{\prime 2}=-\psi_{1}$ from the matrix $g_{\alpha\beta}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ and, similarly, for pointed spinors, $\psi_{\dot{1}}=\psi^{\dot{2}},\psi_{\dot{2}}=-\psi^{\dot{1}}$ [1]. The same author defines bispinors as the pair $(\psi^{\alpha},\varphi_{\dot{\alpha}})$, which form a broader group of Lorentz, and, with them, the scalar product is formulated as $(\psi^{\alpha},\varphi_{\dot{\alpha}})\cdot(f^{\alpha},h_{\dot{\alpha}})$. The author of the reference [9] mentions algebra $Cl(3)$ as a space-time formulation generated by vectors $e_{\mu}$, which form a basis for $\mathds{R}^{1,3}$ that satisfies the relation $e_{\mu}\cdot e_{\nu}=g_{\mu\nu}$, inducing a 16 -membered basis, as described below:
  • 1
    $1$ scalar,
  • 2
    $(e_{0},e_{1},e_{2},e_{3})$ 4-vector,
  • 3
    $(e_{0}e_{1},e_{0}e_{2},e_{0}e_{3},e_{1}e_{2},e_{2}e_{3},e_{3}e_{1})$ 6- bivectors,
  • 4
    $(e_{1}e_{2}e_{3},e_{0}e_{2}e_{3},e_{0}e_{1}e_{3},e_{0}e_{1}e_{2})$ 4-trivectors,
  • 5
    $e_{5}\equiv e_{0}e_{1}e_{2}e_{3}$ pseudoescalar.
In accordance with the above, the same author describes the Weyl spinors as
$\begin{bmatrix}\psi^{1}\\ \psi^{2}\end{bmatrix}$ and the Dirac spinor $\begin{bmatrix}\psi^{\alpha}\\ \varphi^{\dot{\alpha}}\end{bmatrix}$, as defined in the work of Beretetskii et al [1].
The $q$ Lorentzian algebra was defined in the reference [12]. The quantum complex spinors have components $\psi_{1}$ and $\psi^{2}$ and conjugates $\varphi^{\dot{1}}$ and $\varphi_{\dot{2}}$. For all $q\in\mathds{R}-\left\{0\right\}$, they satisfy the following $q$ relations
$\psi_{1}\psi^{2}=q\psi^{2}\psi_{1},$ $\ \psi^{2}\varphi^{\dot{1}}=q\varphi^{\dot{1}}\psi^{2},$ (2)
$\psi_{1}\varphi^{\dot{1}}=\varphi^{\dot{1}}\psi_{1}-q(q+q^{-1})^{1/2}\varphi_{\dot{2}}\varphi^{\dot{1}},$ $\ \psi^{2}\varphi^{\dot{1}}=\varphi^{\dot{1}}\psi^{2},$ (3)
$\psi_{1}\varphi_{\dot{2}}=q\varphi_{\dot{2}}\psi_{1},$ $\ \varphi^{\dot{1}}\varphi_{\dot{2}}=q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}}.$ (4)
Considering spinors $\psi^{\alpha},\varphi^{\dot{\alpha}}$ and $\left\{\tau^{1},T^{2},S^{1},\sigma^{2}\right\}$ as the generators of the $q$ Lorentzian algebra for the group $U_{q}(su(2))$ [13], we have:
  • 1.
    For $\psi^{\alpha},\alpha=1,2$
    $\tau^{1}\psi_{1}=$ $\ \psi_{1}\tau^{1},$ (5)
    $\tau^{1}\psi^{2}=$ $\ \psi^{2}\tau^{1}-q(q+q^{-1})^{2}\psi_{1}T^{2},$ (6)
    $T^{2}\psi_{1}=$ $\ q^{-1}\psi_{1}T^{2},$ (7)
    $S^{1}\psi_{1}=$ $\ q\psi_{1}S^{1},$ (8)
    $T^{2}\psi^{2}=$ $\ q\psi^{2}T^{2},$ (9)
    $S^{1}\psi^{2}=$ $\ q^{-1}\psi^{2}S^{1}-\psi_{1}\sigma^{2},$ (10)
    $\sigma^{2}\psi_{1}=$ $\ \psi_{1}\sigma^{2},$ (11)
    $\sigma^{2}\psi^{2}=$ $\ \psi^{2}\sigma^{2},$ (12)
  • 2.
    Their complex conjugates $\varphi^{\dot{\alpha}},\alpha=\dot{1},\dot{2}$
    $\tau^{1}\varphi^{\dot{1}}=$ $\ q^{-1}\varphi^{\dot{1}}\tau^{1},$ (13)
    $\tau^{1}\varphi_{\dot{2}}=$ $\ q\varphi_{\dot{2}}\tau^{1},$ (14)
    $T^{2}\varphi^{\dot{1}}=$ $\ \varphi^{\dot{1}}T^{2}+q^{-1}\varphi_{\dot{2}}\tau^{1},$ (15)
    $T^{2}\varphi_{\dot{2}}=$ $\ \varphi_{\dot{2}}T^{2},$ (16)
    $S^{1}\varphi^{\dot{1}}=$ $\ \varphi^{\dot{1}}S^{1},$ (17)
    $\sigma^{2}\varphi^{\dot{1}}=$ $\ q\varphi^{\dot{1}}\sigma^{2}+(q+q^{-1})^{2}\varphi_{\dot{2}}S^{1},$ (18)
    $\sigma^{2}\varphi_{\dot{2}}=$ $\ q\varphi_{\dot{2}}\sigma^{2}.$ (19)
Deformed commutation relations for $q$- Lorentzian algebra are defined in the next proposition, on the quantum-symmetric plane and the quantum anti-symmetric plane.
Proposition 1.1.
Consider generator $T$ of the set $\left\{\tau^{1},T^{2},S^{1},\sigma^{2}\right\}$ for the algebra $U_{q}(su(2))$ and the relations 2, 3, and 4 , defined in [12] and [13] . The $q$ Lorentzian algebra for spinors in the deformed space is defined through the following relations:
$T(\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})=$ $\ (\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})T,$ (20)
$T(\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})=$ $\ (\psi_{1}\psi^{2}-q\psi{2}\psi^{1})T,$ (21)
$T(\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})=$ $\ (\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})T,$ (22)
$T(\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})=$ $\ (\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})T.$ (23)
Definition 1.2.
The following are the bosonic $q$ deformed Minkowskian Pauli spin matrices defined in the Schmidt work [13]:
$(\sigma^{+})_{\alpha\dot{\beta}}=\begin{bmatrix}0&0\\ 0&q\end{bmatrix},\quad(\sigma^{-})_{\alpha\dot{\beta}}=\begin{bmatrix}q&0\\ 0&0\end{bmatrix},$ (24)
$(\sigma^{3})_{\alpha\dot{\beta}}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\sigma^{0})_{\alpha\dot{\beta}}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (25)
Likewise, the conjugated Pauli matrices are:
$(\overline{\sigma}^{+})_{\dot{\alpha}\beta}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\overline{\sigma}^{-})_{\dot{\alpha}\beta}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (26)
$(\overline{\sigma}^{3})_{\dot{\alpha}\beta}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\overline{\sigma}^{0})_{\dot{\alpha}\beta}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{-1/2}\\ -q^{1/2}&0\end{bmatrix}.$ (27)
The inverse Pauli matrices
$(\sigma^{-1}_{+})_{\alpha\dot{\beta}}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\sigma^{-1}_{-})_{\alpha\dot{\beta}}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (28)
$(\sigma^{-1}_{3})_{\alpha\dot{\beta}}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\sigma^{-1}_{0})_{\alpha\dot{\beta}}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (29)
Finally
$(\overline{\sigma}^{-1}_{+})_{\dot{\alpha}\beta}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\overline{\sigma}^{-1}_{-})_{\dot{\alpha}\beta}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (30)
$(\overline{\sigma}^{-1}_{3})_{\dot{\alpha}\beta}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\overline{\sigma}^{-1}_{0})_{\dot{\alpha}\beta}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (31)
Definition 1.3.
The $q$ Lorentzian spinor variables or $q$ spinor variables are defined according to the expressions (2), (3), and (4) as follows:
$u^{2}_{1}\equiv$ $\ \psi_{1}\psi^{2}-q\psi^{2}\psi_{1},$ (32)
$v^{\dot{1}2}\equiv$ $\ \psi^{2}\varphi^{\dot{1}}-q\varphi^{\dot{1}}\psi^{2},$ (33)
$x^{\dot{1}}_{1\dot{2}}\equiv$ $\ \psi_{1}\varphi^{\dot{1}}-\varphi^{\dot{1}}\psi_{1}+q(q+1)^{-1/2}\varphi_{\dot{2}}\varphi^{\dot{1}},$ (34)
$y^{2\dot{1}}\equiv$ $\ \psi^{2}\varphi^{\dot{1}}-\varphi^{\dot{1}}\psi^{2},$ (35)
$z^{\dot{1}}_{\dot{2}}\equiv$ $\ \psi_{1}\varphi_{\dot{2}}-q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}},$ (36)
$t_{1\dot{2}}\equiv$ $\ \psi_{1}\varphi_{\dot{2}}-q\varphi_{\dot{2}}\psi_{1}.$ (37)
Definition 1.4.
We consider the set $U=\left\{u^{2}_{1},v^{\dot{1}2},x^{\dot{1}}_{1\dot{2}},z^{\dot{1}}_{\dot{2}},y^{2\dot{1}},t_{1\dot{2}}\right\}\subset\mathds{C}$. A function on the $q$ spinor variables is defined as $\Psi(U)=\Psi(u^{2}_{1},v^{\dot{1}2},x^{\dot{1}}_{1\dot{2}},z^{\dot{1}}_{\dot{2}},y^{2\dot{1}},t_{1\dot{2}})$.
Definition 1.5.
Let $f,g:U\longrightarrow\mathds{C}$ be functions and $u^{\beta}\in U$. The following properties are satisfy on the $q$ spinor variables, we state some clear properties of the functions on the $q$ spinor variables
  • 1.
    $(f+g)(u^{\beta})=f(u^{\beta})+g(u^{\beta})$.
  • 2.
    $(f\cdot g)(u^{\beta})=f(u^{\beta})\cdot g(u^{\beta})$.
  • 3.
    $(f-g)(u^{\beta})=f(u^{\beta})-g(u^{\beta})$.
  • 4.
    $\left(\dfrac{f}{g}\right)(u^{\beta})=\dfrac{f(u^{\beta})}{g(u^{\beta})},\quad g(u^{\beta})\neq 0$.
Definition 1.6.
For a function $f:U\longrightarrow\mathds{C}$ and $u^{\beta}\in\mathbb{C}$ , the $q$ spinor derivative is defined as [8] :
$\dfrac{\mathrm{d}_{q}f}{\mathrm{d}_{q}u^{\beta}}=\dfrac{f((qu)^{\beta})-qf(u^{\beta})}{(qu)^{\beta}-qu^{\beta}},$ (38)
and its conjugate complex
$\dfrac{\mathrm{d}_{q}f}{\mathrm{d}_{q}v^{\dot{\alpha}}}=\dfrac{f((qv)^{\dot{\alpha}})-qf(v^{\dot{\alpha}})}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}}.$ (39)
1.1 Clifford algebra and Dirac operator
Let $\left\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n},\right\}$ be an orthonormal basis of $\mathds{R}^{n}$. The Clifford algebra is generated over $\mathbb{R}^{n}$ under the relation
$\gamma_{\mu}\gamma_{\nu}+\gamma_{\nu}\gamma_{\mu}=-2\delta_{\mu\nu}\gamma_{0},\quad\gamma_{\mu}^{2}=-|\gamma_{\mu}|^{2}\gamma_{0},\quad\mu,\nu=1,2,...,n,$ (40)
where $\delta_{\mu\nu}$ is the Kronecker symbol (see [11], [3], [10] for more details). We will denote the Clifford algebra by $Cl_{n}$, and each element in $Cl_{n}$ can be expressed by its components as $\sum_{a}\gamma_{a}x_{a}$, where $a=(\mu_{1},...,\mu_{n})$ with each $\mu_{l}\in\left\{1,2,...,n\right\}$. Any element $\textit{{x}}\in\mathds{R}^{n}$ can be identified with a 1-vector in the Clifford algebra [10]
$(x_{1},x_{2},...,x_{n})\longrightarrow\textit{{x}}=x_{1}\gamma_{1}+x_{2}\gamma_{2}+\cdots+x_{n}\gamma_{n}.$ (41)
On other hand, the Dirac operator used here is
$D:=\sum\limits_{\mu=1}^{n}\gamma_{\mu}\frac{\partial}{\partial x_{\mu}},$ (42)
we refer to reader to [3], [6], [4] for more details.
1.2 $q$-Deformed Dirac matrices
Definition 1.7.
The $q$-deformed Dirac matrices are defined in [15], and are given by
$\gamma_{\mu}:=\ \begin{bmatrix}0&(\sigma_{\mu})^{\alpha}_{\dot{\beta}}\\ (\overline{\sigma}_{\mu})^{\dot{\alpha}}_{\beta}&0\end{bmatrix},$ (43)
where $(\sigma_{\mu})^{\alpha}_{\dot{\beta}}$ and $(\overline{\sigma}_{\mu})^{\dot{\alpha}}_{\beta}$ denote the Pauli matrices of $q-$ deformed Minkowski space (e.g. [14] for more details), and are defined as
$\begin{split}(\sigma_{+})^{\alpha}_{\dot{\beta}}=\begin{bmatrix}0&0\\ 0&kq^{1/2}\lambda^{1/2}_{+}\end{bmatrix},\quad(\sigma_{3})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}0&q\\ 1&0\end{bmatrix},\\ (\sigma_{-})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}q^{1/2}\lambda^{1/2}_{+}&0\\ 0&0\end{bmatrix},\quad(\sigma_{0})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}0&-q^{-1}\\ 1&0\end{bmatrix},\end{split}$ (44)
and their conjugated counterparts
$\begin{split}(\overline{\sigma}_{+})^{\alpha}_{\dot{\beta}}=\begin{bmatrix}0&0\\ 0&\overline{k}q^{-1/2}\lambda^{1/2}_{+}\end{bmatrix},\quad(\overline{\sigma}_{3})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}0&1\\ q^{-1}&0\end{bmatrix},\\ (\overline{\sigma}_{-})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}q^{-1/2}\lambda^{1/2}_{+}&0\\ 0&0\end{bmatrix},\quad(\overline{\sigma}_{0})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}0&1\\ -q&0\end{bmatrix},\end{split}$ (45)
where $k,\overline{k}$ are characteristic parameters associated to bosons ($q=+1$) and fermions ($q=-1$).
1.3 $q$-Spinor complex integral formulas
Definition 1.8.
[8] Let $\Gamma_{q}$ be the closed contour of the deformed quantum complex plane, and $u_{0}^{\beta},v^{\dot{\alpha}}\subset\Gamma_{q}$ point spinors contained in the contour. The $q$- spinor complex integral formulas are defined by
$\oint_{\Gamma_{q}}\dfrac{\Psi(u^{\beta})\mathrm{d}_{q}u^{\beta}}{(qu)^{\beta}-qu^{\beta}_{0}}$ $=\ \dfrac{1}{q}\sum_{n=0}^{\infty}\left[(\overline{\sigma}_{\mu})_{\dot{\alpha}\beta}\Psi(u_{0}^{\beta})\right]^{n},$ (46)
$\oint_{\Gamma_{q}}\dfrac{\Psi((qu)^{\beta})\mathrm{d}_{q}u^{\beta}}{(qu)^{\beta}-qu^{\beta}_{0}}$ $=\ \sum_{n=0}^{\infty}\left[(\overline{\sigma}_{\mu})_{\dot{\alpha}\beta}\Psi((qu_{0})^{\beta})\right]^{n},$ (47)
$\oint_{\Gamma_{q}}\dfrac{\Psi(v^{\dot{\alpha}})\mathrm{d}_{q}v^{\dot{\alpha}}}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}_{0}}$ $=\ \dfrac{1}{q}\sum_{m=0}^{\infty}\left[(\sigma_{\mu})_{\alpha\dot{\beta}}\Psi(v_{0}^{\dot{\alpha}})\right]^{m},$ (48)
$\oint_{\Gamma_{q}}\dfrac{\Psi((qv)^{\dot{\alpha}})\mathrm{d}_{q}v^{\dot{\alpha}}}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}_{0}}$ $=\ \sum_{m=0}^{\infty}\left[(\sigma_{\mu})_{\alpha\dot{\beta}}\Psi((qv_{0})^{\dot{\alpha}})\right]^{m}.$ (49)
Motivation
The topic of this article is $q$-differential operators for $q$-spinor variables, the motivation comes from the study of the differential and integral calculus in $q$-spinor variables studied in [8] and some differential operator, in particular the Dirac operator mentioned in references [3], [4], [6]. In accordance with the above, our interest here is to relate the $q$-spinor variables with some differential operators, especially the $q$-Dirac operator, and its implications with differential calculus, integral and differential equations. The main objective of this work is, therefore, to study the $q$-differential operator on the $q$-spinor variables and its corresponding differential and integral calculus, and its differential equations. Solutions to the differential equation in $q$-spinor variables were also found.
This paper is organized as follows. We briefly recall the preliminaries will be used in this paper in Sect.2. The $q$-differential operators for $q$-spinor variables, the $q$-spinor chain rule, the new $q$-differential operator, the $q$-Dirac differential operator, and the integral formulas in $q$-spinor variables are then proposed in Sect. 3. Finally in the last Section the discussion and some suggestions for further work are presented.
Notation
Taking into account the Definition 1.5 the functions on the $q$ spinor variables we will denote by $\Psi(u^{\alpha}_{\dot{\beta}})$ for all $\alpha,\dot{\beta}=1,2$ , and thus we can write (38) and (39) of the form
$\dfrac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}=\dfrac{\Psi((qu)^{\alpha}_{\dot{\beta}})-q\Psi(u^{\alpha}_{\dot{\beta}})}{(qu)^{\alpha}_{\dot{\beta}}-qu^{\alpha}_{\dot{\beta}}},\quad(qu)^{\alpha}_{\dot{\beta}}\neq qu^{\alpha}_{\dot{\beta}}.$ (50)

### v2

@@ Abstract
We introduce a q-differential operator adapted to q-spinor variables, establishing a corresponding q-spinor chain rule and defining both standard and Dirac-type q-differential operators. Integral formulas in q-spinor variables are derived, and applications to q-deformed spinor differential equations are explored through explicit examples. The framework extends existing q-calculus to spinorial structures, offering potential insights into quantum deformations of relativistic field equations. We conclude with suggestions for future developments, including a q-analogue of the Dirac–Maxwell algebra.
@@ 1 Introduction
The mathematical formulation of spinor fields is essential for the study of relativistic quantum mechanics, with foundations established in the works of Cartan [2], Berestetskii et al. [1], and others. Spinors are typically represented as two-component objects, such as $\psi^{\alpha}=\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}$, and their conjugates are associated with transformation properties under the Lorentz group. The representation theory underlying these structures connects to the group $SU(2)$, whose generators are the Pauli matrices [17].
Gori et al. [5] provide explicit matrix forms for rotations, revealing the origin of the Pauli matrices from representations of spinor transformations. Berestetskii et al. [1] formalised the covariant and contravariant behavior of spinors, as well as their scalar products and bispinor structures. Lachieze-Rey [10] further connected these ideas to Clifford algebras, particularly $Cl(3)$, which provides a geometric interpretation of space-time through multivector bases in $\mathds{R}^{1,3}$.
In the context of quantum deformations, q-spinor variables and their associated algebraic relations emerge from the q-deformed Lorentz algebra [12]. These variables satisfy non-trivial commutation relations, forming the foundation for defining q-differential operators. The structure of the q-Lorentzian algebra and its generators, such as $\tau^{1},T^{2},S^{1},\sigma^{2}$, governs the algebraic relations between spinors and their conjugates [13].
Building on the formalism developed in [8], we define q-spinor variables and derive corresponding differential and integral operators, including q-derivatives and q-complex integrals. These operators act on functions defined over quantum spinor variables and respect the q-deformed algebraic relations. Moreover, the formulation introduces q-Pauli matrices and q-Dirac matrices [15], extending classical Clifford structures to a deformed setting.
We also review the classical Dirac operator defined over Clifford algebras [11, 3, 4], and reinterpret its structure in terms of q-deformed analogues. The resulting q-Dirac operator inherits many key properties of its classical counterpart, including its role in defining and solving differential equations.
The principal motivation of this work is to investigate the role of q-differential operators in spinor calculus, as introduced in [8], and to explore their potential in the formulation and solution of q-deformed differential equations. Inspired by earlier treatments of classical and quantum Dirac operators [6], our objective is to construct a consistent differential and integral calculus over q-spinor variables, develop explicit solutions to q-Dirac-type equations, and outline the implications of this theory for quantum-deformed field equations.
We conclude the introduction by highlighting the long-term goal of formulating a q-Dirac–Maxwell algebra and a q-real spinor calculus, drawing from both geometric and algebraic perspectives [16]. These efforts aim to enrich the interplay between quantum groups, spin geometry, and field theory, laying the groundwork for future developments in noncommutative quantum physics. This paper is organized as follows. We briefly recall the preliminaries will be used in this paper in Section 2. The $q$-differential operators for $q$-spinor variables, the $q$-spinor chain rule, the new $q$-differential operator, the $q$-Dirac differential operator, and the integral formulas in $q$-spinor variables are then proposed in Section 3. Finally in the last Section the discussion and some suggestions for further work are presented.